OpenStudy (2yoututu11):

Solve. 4y = 144 A. 36 B. 16 C. –16 D. –36

OpenStudy (anonymous):

Divide 4 on both sides

OpenStudy (anonymous):

y=144/4 38.5

OpenStudy (anonymous):

4y/4 = 144/4 y=36